# Composed Multi-modal Retrieval: A Survey of Approaches and Applications

### v1

@@ 3.2.2 Design of Plug and Play Structure
Based on the key highlights of different approaches in achieving plug-and-play structures, we classify them into the following four categories: (1) Modular design to enhance plug-and-play capability [54, 55]; (2) Dynamic configuration and adaptive mechanisms [56, 57]; (3) Data alignment and feature fusion techniques [28, 46]; (4) Training strategies and re-ranking improvements. Below, we introduce the details of each category [50, 4].
(1) Modular design to enhance plug-and-play capability: To improve the flexibility and scalability of multi-modal retrieval systems, some methods adopt modular network architecture designs, allowing these solutions to be easily integrated into existing systems. Chen et al. [54] proposed a Joint Visual Semantic Matching (JVSM) model that associates visual and textual modalities by constructing a shared discriminative embedding space, enabling image-text matching. The JVSM employs a four-module architecture (visual embedding, text embedding, semantic projection, and combination module), optimized with various loss functions, allowing it to be seamlessly integrated into existing systems without major architectural changes. Experimental results demonstrate its superior performance in handling complex queries, showcasing strong plug-and-play capabilities. Yang et al. [55] focused on addressing asymmetries and heterogeneities between cross-modal data, proposing a novel alignment loss function to support effective feature mapping and alignment. This method is designed as an independent component that can be directly applied to existing CMR systems, enhancing the association between different modalities and improving system flexibility and adaptability.
(2) Dynamic configuration and adaptive mechanisms: Some works emphasize dynamic adjustment capabilities based on specific application scenarios, allowing systems to flexibly respond to different query types and user needs. Kim et al. [56] developed a dual compositional learning framework capable of dynamically adjusting composition and fusion strategies according to specific application scenarios, adapting to changes in user queries, thus significantly enhancing interactive image retrieval performance. This flexible design enables systems to quickly adapt to new environments or requirements, demonstrating high flexibility and ease of integration. Zhu et al. [57] introduced an adaptive multi-expert collaborative network that selects the most suitable expert models for processing inputs, providing more personalized and precise retrieval services. This dynamic selection mechanism based on input characteristics allows to rapidly adapt and integrate into various text-guided image retrieval systems, greatly enhancing system functionality and robustness.
(3) Data alignment and feature fusion techniques: Addressing asymmetries and heterogeneities between heterogeneous data, researchers have proposed various feature alignment and fusion techniques to enhance cross-modal retrieval performance. Pang et al. [28] presented a framework that combines multiple types of features and ensures accurate correspondence through cross-modal alignment algorithms. The authors designed a multi-layer perceptron for fusing heterogeneous features and used contrastive loss for aligning different modalities. This method not only improves the relevance and accuracy of retrieval results but also demonstrates its potential as a plug-and-play component. Zhang et al. [46] proposed a comprehensive relationship reasoning mechanism that analyzes relationships between images and texts in detail, constructing a refined relational graph and using graph convolutional networks to capture these relationships. This comprehensive reasoning approach can be directly applied to existing CMR systems, significantly enhancing the performance of composed query-based image retrieval.
(4) Training strategies and re-ranking improvements: Finally, several studies explored how to directly enhance retrieval system performance through improved training strategies and candidate set re-ranking. Liu et al. [50] utilized dual multi-modal encoders for the candidate set re-ranking, offering a convenient way to optimize final retrieval results. This lightweight design allows for rapid implementation without altering the original system architecture, thereby improving system efficiency and accuracy. Liu et al. [4] investigated a bi-directional training strategy via text prompt learning, emphasizing the importance of lightweight design, enabling quick deployment without changing the original system architecture, and further enhancing system understanding and retrieval capabilities. This highlights the advantages of bi-directional training in improving retrieval effectiveness and demonstrates its efficacy as a plug-and-play component.
3.2.3 Design of Non-combiner Architecture
@@
Non-combinatorial structural design refers to methods that avoid directly merging or combining different modalities of data (such as text and images) when processing multimodal data. Instead, these approaches achieve cross-modal information fusion and utilization through indirect means, aiming to circumvent the complexity and limitations that traditional combinatorial methods may introduce. Specifically, non-combinatorial structural design attempts to preserve the uniqueness of each modality’s data through various strategies, and on this basis, conduct information interaction, enhancement, and complementation to achieve more effective information retrieval and utilization.
For instance, Zhang et al.[8] proposed the Tell-Imagine-Search framework, which consists of three modules. The Tell module generates a detailed text description based on the input text and reference image; the Imagine module synthesizes an imaginary image according to this description; and the Search module uses the synthesized image to search for the most similar real image in the database. This indirect cross-modal information fusion approach improves retrieval accuracy and enhances the system’s creativity and flexibility. Subsequently, Delmas et al.[58] proposed the ARTEMIS framework, which utilizes attention mechanisms to enhance cross-modal retrieval performance. ARTEMIS allows the model to simultaneously consider explicit matching of text content and implicit similarities between image features, introduces a new loss function to guide learning, and adopts a lightweight network architecture to ensure inference efficiency, reflecting the preservation and utilization of each modality’s uniqueness.
Fu et al.[59] constructed a Multi-Order Adversarial Representation Learning (MOARL) method to address representation learning. MOARL includes pixel-level, feature-level, and semantic-level multi-order adversarial training, balancing the learning process at different levels through specially designed loss functions, showcasing its non-combinatorial structural characteristics and improving the model’s understanding of complex queries. Han et al.[60] proposed the FAME-ViL multi-task model, which employs a shared encoder-decoder structure to tackle the challenges of heterogeneous vision-language tasks in the fashion domain. The encoder extracts universal vision-language features, while different decoders correspond to specific tasks such as product classification, attribute prediction, and image retrieval. This architecture, which processes multimodal information in stages, demonstrates the flexibility and efficiency of non-combinatorial structural design. Yang et al.[61] proposed the Decompose Semantic Shifts (FSS) method to address the semantic shift problem. FSS identifies key query elements that may cause semantic shifts through semantic parsing and learns stable representations of these elements through a contrastive learning framework for precise image retrieval. This showcases the potential of non-combinatorial structural design in enhancing the retrieval relevance and robustness of complex queries. Wen et al.[62] transferd multimodal fusion to the raw data level, leveraging vision-language pre-training models to achieve multimodal encoding and cross-modal alignment. This framework resolves the embedding space shift problem that may arise from feature-level fusion, further validating the effectiveness of non-combinatorial structural design.
@@ 3.3.1 Based on Contrastive Learning
$L=[\kappa(\psi_{i},\phi^{+}_{i})-\kappa(\psi_{i},\phi^{-}_{i,j})+\gamma]_{+}$ (7)
@@
In addition to designing the loss function, existing work also focuses on the negative sample mining strategies and the similarity metric function.
Negative Sample Mining: As a fundamental role in training, there are two common strategies: mini-batch hardest negative mining [49, 65, 66] and mini-batch semi-hard negative mining [54, 33]. The former only selects negative examples with the greatest similarity to the combined query features, while the latter selects negative examples with a higher similarity to the combined query than the target image. In Work A, it is validated that semi-hard mining is more stable and faster converging than hardest mining.
Unlike simple selection, Zhang et al. [3] propose a method to enhance multimodal fusion by replacing parts of the query’s text to generate harder negative samples, improving model learning. They introduce two techniques for mining hard examples: (1) synthetic composition examples, which create negative samples by replacing the text while keeping the image part unchanged, forming a new sample as $f=r+t^{\prime}$, where $t^{\prime}$ is from another query. (2) augmented synthetic composition examples, which replace only parts of the text, generating even harder-to-distinguish negatives. A mask vector $s$ randomly determines which dimensions of the text to retain or replace, with a Bernoulli distribution governing the decision. The new text embedding is $\tilde{t}^{\prime}=s\odot t+(1-s)\odot t^{\prime}$. They also introduce a dynamic replacement ratio, sampling $p$ from a Beta distribution for more diverse hard examples.
@@
$L=\frac{1}{B}\sum_{i=1}^{B}-\log\frac{\exp(\kappa(\psi_{i},\phi^{+}_{i}))}{\exp(\kappa(\psi_{i},\phi^{+}_{i}))+\sum_{j=1}^{B-1}\exp(\kappa(\phi^{+}_{i},f_{j}))}$ (10)
@@
Similarity Metric Function: Conventional contrastive learning losses typically employ Euclidean distance or cosine similarity as metrics for measuring the distance between samples [18, 12, 10]. However, these methods consider only the absolute distance or angular relationship between samples, neglecting the combined effect of both factors. To address this limitation, Zhang et al. [29] propose the Triangle Area (TA) as a novel metric for sample distance measurement. Specifically, as illustrated in Figure A, for an anchor sample $a$, a positive sample $c$, and a negative sample $b$, the distances between the samples are represented by the areas of triangles $Oab$ and $Oac$, where $O$ denotes the origin. TA is defined as:
$\text{TA}(\psi_{i},\varphi_{i})=\frac{1}{2}|\psi_{i}||\phi_{i}|\sqrt{1-\left(\frac{\psi_{i}\cdot\phi_{i}}{|\psi_{i}||\phi_{i}|}\right)^{2}}$ (11)
TA offers two main advantages: (1) It considers both the absolute distance and angular relationship between samples, allowing the model to optimize both aspects without needing manual weight adjustments. (2) For smaller datasets, TA uses a squared distance, amplifying differences and slowing convergence to prevent overfitting. For larger datasets, it uses the area as a distance metric, avoiding excessive magnification of differences and enabling faster convergence and reduced training time.
@@ 3.3.2 Consistency Constraint
Works [35, 30, 23] focus on consistency loss for image and text reconstruction, aiming to construct a robust and meaningful joint representation space. Within this space, the model attempts to reconstruct the original input images and texts as faithfully as possible, learning their associations and semantic relationships to improve retrieval accuracy. Additionally, works [41] and [36] delve into the reconstruction consistency between global and local composite features. For instance, Wen et al.[41] utilize fine-grained local and global combination modules to achieve multimodal integration and introduce an inter-enhancement module to promote mutual improvement between local and global features. It enhances knowledge transfer across different feature levels using KL divergence loss and feature-level $L_{2}$-norm loss to optimize the consistency between local and global features. This approach not only improves the expressiveness of local features but also reinforces the synergy between the global and local combination modules.
Pang et al. and Zhang et al. [28, 27] introduce Relative Caption-aware Consistency (RCC) loss to address the heterogeneous semantic gap between images and texts. RCC promotes effective alignment and cross-modal fusion by minimizing the discrepancy between generated text snippets and original text snippets. Specifically, RCC employs an improved LSTM structure to generate, word by word, a textual description of the differences between reference and target images. It uses cross-entropy loss to measure the gap between the probability distribution of the generated text and the true labels. By minimizing this loss, the model adjusts its parameters during training to make the generated text closer to the user-provided modified text, thus achieving effective alignment of different modal features in a unified embedding space.
From the perspective of consistency constraints between multi-branch networks, works [13, 12, 61, 57] propose various mechanisms to ensure that different branches or expert networks produce consistent feature representations when processing the same input, thereby enhancing the model’s robustness and generalization capability. Zhang et al. [13] introduce the consensus network, which uses multiple diverse compositors to generate image-text embeddings, each providing a unique perspective. They incorporate KL consistency constraints through Kullback-Leibler divergence loss to promote mutual learning among different compositors, reducing individual biases and reaching consensus, which increases robustness against noisy annotations. Wen et al. [12] design a teacher-student framework with two branches: the teacher branch guides directly based on target-query relationships, while the student branch mimics the behavior of the teacher. They use KL consistency constraints to ensure that the student branch learns consistent conflict modeling from the teacher branch, reducing bias and improving the effectiveness of multimodal query combinations and candidate image ranking quality.
Yang et al. [61] propose a Semantic Shift Network (SSN), which decomposes text instructions into degradation and upgradation steps, explicitly handling the transition from reference images to target images. Through KL consistency constraints, SSN ensures that the positive sample feature distribution aligns with the target distribution while increasing the difference between negative samples and the target image, enhancing the accuracy of positive sample features and the model’s ability to distinguish between positive and negative samples. Zhu et al. [57] present an adaptive multi-expert collaborative network, which introduces multiple expert networks focusing on specific types of feature extraction, such as local details, global structures, or semantic relationships. They implement consistency constraints to minimize the differences in feature representations among different expert networks, encouraging similar outputs when processing the same input, thereby enhancing overall performance and robustness. The core idea of these methods is to design appropriate consistency constraints so that multi-branch or multi-expert networks can learn from each other during training, reducing their respective biases.
@@ 4 Zero-Shot Learning-based CMR
Without relying on triplet annotations, thereby avoiding limitations caused by data, existing methods focus on simulating compositional retrieval based solely on image-text information, establishing a zero-shot learning mechanism. Popular paradigms can be divided into three categories: (1) using images as input, designing a learning framework for visual-to-language transformation to combine modified text with query target images at test time; (2) using text as input, simulating the visual-to-language transformation process, and replacing text input with visual image input at test time, which offers higher training efficiency; (3) introducing additional knowledge for enhancement. The following sections will introduce the progress in each of these areas.
@@ 4.1 Image-side Transformation
In existing zero-shot compositional multi-modal retrieval, the process of image-to-visual transformation can be categorized into explicit and implicit approaches. The explicit approach directly converts input images into explicit textual descriptions, typically using pre-trained captioning models due to their training-free nature, but it is limited by the performance of these captioning models. The implicit approach, on the other hand, transforms input images into implicit feature vectors, usually encoded in a word vector space. This method is learnable and more flexible, allowing for a more sophisticated design tailored to the specific characteristics of the task.
@@ 4.1.1 Explicit visual Transformation (Training-free)
Existing explicit visual transformation methods [67, 68, 69] utilize pre-trained visual-language models (VLMs) and large language models (LLMs) to achieve zero-shot composed image retrieval (ZS-CIR) in a training-free manner.
Karthik et al. [67] propose a modular framework CIReVL, which generates captions for reference images and uses LLMs to recompose these captions based on the textual target modifications. This approach avoids the need for additional adaptation resources, supports flexible model component replacements, and allows user-level interventions by refining captions post-hoc, achieving competitive or superior results on four CIR benchmarks. Sun et al. [68] introduce a two-stage ranking method to address modality gaps and ambiguous requirements from the reference image. The first stage converts composed image-text queries into text-only queries for global retrieval, while the second stage re-ranks top-K results by extracting and evaluating local attributes from the modified instructions. This method combines global and local information to outperform other training-free approaches on open-domain datasets. Yang et al. [69] develop LDRE, which employs dense caption generation for reference images and uses LLMs for divergent reasoning to create a range of potential target image captions. An ensemble strategy integrates these candidates to better align with the query intent, effectively handling complex or ambiguous queries. Despite their differences in visual transformation strategies, all methods rely on pre-trained models and language-driven reasoning to bridge the gap between modalities and enable effective, training-free composed retrieval.

### v2

@@ 3.2.2 Design of Plug and Play Structure
Based on the key highlights of different approaches in achieving plug-and-play structures, we classify them into the following two categories: (1) modular design to enhance plug-and-play capability [69, 128, 66, 59]; and (2) Training and inference strategies with plug-and-play characteristics [93, 92, 43, 215].
Focusing on devising plug-and-play modules to enhance system flexibility and performance, Chen et al. [69] propose a Joint Visual Semantic Matching (JVSM) module that associates visual and textual information through a shared discriminative embedding space. Their four-module architecture, encompassing visual embedding, textual embedding, semantic projection, and composition modules, seamlessly integrates into existing retrieval systems, effectively handling image-text matching tasks, particularly excelling in complex fashion term retrieval. Meanwhile, Kim et al. [128] introduce a dual compositional module, which incorporates a correction network to complement traditional compositional networks by capturing differences between reference and target images. This dual mapping design not only generates composite features but also models image discrepancies, allowing it to be flexibly integrated into diverse retrieval systems. Yang et al. [68] propose a cross-modal joint prediction module, in which the modified text is treated as an implicit transformation between the query image and the target image. This module can be seamlessly integrated into existing methods to enhance the discriminability and robustness of visual and textual representations.
Besides, Zhu et al. [66] present an adaptive multi-expert collaborative module, integrating multiple expert models focused on specific image regions or text segments, with an adaptive gating mechanism dynamically adjusting weights based on input. This approach facilitates more refined feature extraction and composition, improving the system’s capability to handle complex and varied queries, demonstrating strong applicability as a plug-and-play component. Zhang et al. [59] develop a comprehensive relationship reasoning module, utilizing graph convolutional networks to analyze multiple relationships between images and texts, such as contextual and category relationships. This module improves the relevance and accuracy of retrieval results.
Focusing on plug-and-play design in training and inference strategies, Liu et al. [93] employ a bidirectional training strategy to optimize cross-modal interactions, improving model adaptability and overall retrieval accuracy, especially in few-shot scenarios. Hou et al. [92] construct pseudo-triplets using anchor samples to simulate semantic relationships and generate discriminative training instances, effectively alleviating data scarcity and demonstrating its potential as a lightweight plug-and-play component. Pang et al. [43] address heterogeneous feature fusion and cross-modal alignment by integrating multiple visual attributes via an MLP and applying contrastive loss to unify cross-modal semantics, enhancing the model’s ability to understand complex compositional queries while maintaining modularity. Liu et al. [215] present a dual multi-modal encoder for candidate set re-ranking in the inference stage, improving relevance and ranking quality with strong integration capabilities into retrieval systems.
3.2.3 Design of Non-explicit Combiner Architecture
@@
Non-combinatorial structural design refers to methods that avoid directly fusing different modalities of data (such as images and texts) when processing multimodal data. Instead, these approaches achieve cross-modal information fusion and utilization through indirect means, aiming to circumvent the complexity and limitations that traditional combinatorial methods may introduce.
Specifically, non-combinatorial structural design attempts to preserve the uniqueness of each modality’s data through various strategies, and on this basis, conducts information interaction, enhancement, and complementation to achieve more effective information retrieval. Zhang et al.[97] propose the Tell-Imagine-Search framework, which consists of three modules. The Tell module generates a detailed text description based on the input text and reference image; the Imagine module synthesizes an imaginary image according to this description; and the Search module uses the synthesized image to search for the most similar real image in the database. Subsequently, Delmas et al.[129] propose a novel approach to split the combined image retrieval task into two independent but complementary subtasks: explicit matching and implicit similarity. Explicit matching uses an attention mechanism to focus on processing visual elements explicitly described in the query text. Implicit similarity considers broader contextual similarities between images.
In addition, Yang et al.[67] propose the decompose semantic shifts (FSS) method, which identifies key query elements that may cause semantic shifts. FSS explicitly decomposes the semantic change into two steps: from reference image to visual prototype and then from visual prototype to target image. This approach not only preserves key visual cues but also enriches the visual prototype with textual guidance. Wen et al.[130] transfer multimodal fusion to the raw data level, leveraging vision-language pre-training models to achieve multimodal encoding and cross-modal alignment. This framework resolves the embedding space shift problem that may arise from feature-level fusion, further validating the effectiveness of non-combinatorial structural design.
@@ 3.3.1 Based on Contrastive Learning
$L=[\kappa(\psi_{i},\phi^{+}_{i})-\kappa(\psi_{i},\phi^{-}_{i,j})+\gamma]_{+},$ (7)
@@
The famous margin-based triplet ranking loss has many advanced variants in cross-modal retrieval. Wang et al. [218] consider the external constraint loss that preserves the neighborhood structure in a single modality. There are also variants such as the ladder loss [132], the polynomial loss [131], and the adaptive offline quintuplet loss [133]. In addition to designing the loss function, existing work also focuses on the negative sample mining strategies and the similarity metric function.
Negative Sample Mining: As a fundamental role in training, there are two common strategies: mini-batch hardest negative mining [62, 219, 70] and mini-batch semi-hard negative mining [69, 48]. The former only selects negative examples with the largest similarity to the combined query features, while the latter selects negative examples with a higher similarity to the combined query than the target image. It is validated that semi-hard mining is more stable and faster converging than the hardest mining [69].
Unlike simple selection, Feng et al. [134] propose a method to expand positive and negative examples through MLLM and design a two-stage fine-tuning framework to optimize the representation space. Zhang et al. [71] propose a method to enhance multimodal fusion by replacing parts of the query’s text to generate harder negative samples, improving model learning. They introduce two techniques for mining hard examples: (1) synthetic composition examples, which create negative samples by replacing the text while keeping the image part unchanged, forming a new sample as $f=r+t^{\prime}$, where $t^{\prime}$ is from another query. (2) augmented synthetic composition examples, which replace only parts of the text, generating even harder-to-distinguish negatives. A mask vector $s$ randomly determines which dimensions of the text to retain or replace, with a Bernoulli distribution governing the decision. The new text embedding is $\tilde{t}^{\prime}=s\odot t+(1-s)\odot t^{\prime}$. They also introduce a dynamic replacement ratio, sampling $p$ from a Beta distribution for more diverse hard examples.
@@
$L=\frac{1}{B}\sum_{i=1}^{B}-\log\frac{\exp(\kappa(\psi_{i},\phi^{+}_{i}))}{\exp(\kappa(\psi_{i},\phi^{+}_{i}))+\sum_{j=1}^{B-1}\exp(\kappa(\phi^{+}_{i},f_{j}))}.$ (10)
@@
Similarity Metric Function: Conventional contrastive learning losses typically employ Euclidean distance or cosine similarity as metrics for measuring the distance between samples [20]. However, these methods consider only the absolute distance or angular relationship between samples, neglecting the combined effect of both factors. To address this limitation, Zhang et al. [44] propose the Triangle Area (TA) as a novel metric for sample distance measurement. Specifically, as illustrated in Figure A, for an anchor sample $a$, a positive sample $c$, and a negative sample $b$, the distances between the samples are represented by the areas of triangles $Oab$ and $Oac$, where $O$ denotes the origin. TA is defined as:
$\text{TA}(\psi_{i},\varphi_{i})=\frac{1}{2}|\psi_{i}||\phi_{i}|\sqrt{1-\left(\frac{\psi_{i}\cdot\phi_{i}}{|\psi_{i}||\phi_{i}|}\right)^{2}}.$ (11)
TA has two key benefits: (1) It considers both the distance and angle between samples, letting the model optimize both without manual weight tuning. (2) For small datasets, it uses squared distance to slow convergence and prevent overfitting; for large datasets, it uses area-based distance to avoid exaggerating differences, leading to faster training.
Beyond TA, researchers have also proposed many novel similarity metric functions in cross-modal alignment, including the feature block weighting similarity [185], the cross-modal feature mapping similarity [188], the cross-modal feature dimensional dependency modeling similarity [220], the conditional similarity based on feature decoupling [221], the dimensionally interpretable similarity [222], and the learnable similarity via enhanced measure-units [223].
@@ 3.3.2 Consistency Constraint
Works [50, 45, 38] focus on consistency loss for image and text reconstruction, aiming to construct a robust and meaningful joint representation space. Within this space, the model attempts to reconstruct the original input images and texts as faithfully as possible, learning their associations and semantic relationships to improve retrieval accuracy. Additionally, works [55] and [51] delve into the reconstruction consistency between global and local composite features. For instance, Wen et al.[55] utilize fine-grained local and global combination modules to achieve multimodal integration and introduce an inter-enhancement module to promote mutual improvement between local and global features. It enhances knowledge transfer across different feature levels using KL divergence loss and feature-level $L_{2}$-norm loss to optimize the consistency between local and global features. This approach not only improves the expressiveness of local features but also reinforces the synergy between the global and local combination modules. Pang et al. [43] and Zhang et al. [42] propose the Relative Caption-aware Consistency (RCC) loss to bridge the semantic gap between image and text modalities. RCC guides cross-modal alignment by minimizing the divergence between generated and ground-truth modification texts.
Several studies [99, 72, 67, 66] introduce consistency constraints across multi-branch or multi-expert networks to enhance model robustness and generalization. Zhang et al. [99] propose a consensus network with diverse combiners generating complementary image-text embeddings, regularized by KL divergence to encourage agreement and reduce annotation noise sensitivity. Wen et al. [72] employ a teacher-student framework, where the student mimics the teacher’s target-query reasoning, guided by KL consistency to improve multimodal query alignment. Yang et al. [67] develop a Semantic Shift Network that decomposes text into upgradation and degradation steps, using KL constraints to align positive samples with targets and increase separation from negatives. Zhu et al. [66] design an adaptive multi-expert network that enforces consistency among expert branches focusing on distinct visual features. Collectively, these approaches leverage consistency losses to promote mutual learning, reduce bias, and improve cross-modal representation.
@@ 4 Zero-Shot Learning-based CMR
Without relying on triplet annotations, existing methods focus on simulating composed retrieval based solely on image-text pairs or unlabeled images, establishing a zero-shot learning mechanism. Popular paradigms can be divided into three categories: (1) using images as input, designing a learning framework for visual-to-language transformation to combine modified text as query to search target images at test time; (2) using text as input, simulating the visual-to-language transformation process, and replacing text input with visual image at test time, which offers higher training efficiency; (3) using linear interpolation of visual and textual features. The following sections will introduce progress in each of these areas.
@@ 4.1 Image-side Transformation
In existing zero-shot composed multi-modal retrieval, as shown in Fig. 4, the process of image-to-visual transformation can be categorized into explicit and implicit approaches. The explicit approach directly converts input images into textual descriptions, typically using pre-trained captioning models due to their training-free nature, but it is limited by the performance of these captioning models. The implicit approach, on the other hand, transforms input images into implicit feature vectors, usually encoded in a word vector space. This method is learnable and more flexible, allowing for a more sophisticated design tailored to the specific characteristics of the task.
Figure 4: Explicit and Implicit Visual Transformation. Sub-figures (a) and (b) are from [74] and [82], respectively.
@@ 4.1.1 Explicit Visual Transformation (Training-Free)
Existing explicit visual transformation methods [74, 75, 76, 137] utilize pre-trained visual-language models (VLMs) and large language models (LLMs) to achieve zero-shot CMR in a training-free manner.
Karthik et al. [74] propose a CIReVL framework, which generates captions for reference images and uses LLMs to recompose these captions based on the textual modifications of the target. This approach avoids the need for additional adaptation resources, supports flexible model component replacements, and allows user-level interventions by refining captions post-hoc, achieving competitive results. Sun et al. [75] introduce a two-stage ranking method to address modality gaps and ambiguous requirements from the reference image. The first stage converts composed image-text queries into text-only queries for global retrieval, while the second stage re-ranks top-K results by extracting and evaluating local attributes from the modified instructions. This method combines global and local information to outperform other training-free approaches on open-domain datasets. Yang et al. [76] develop LDRE, which employs dense caption generation for reference images and uses LLMs for divergent reasoning to create a range of potential target image captions. An ensemble strategy integrates these candidates to better align with the query intent, effectively handling complex or ambiguous queries. Yang et al. [137] propose SEIZE, which generates diverse edited captions for the reference image by combining multiple captions from a pre-trained captioning model with relative text through LLM-based reasoning, enabling training-free ZS-CMR. Despite their differences in visual transformation strategies, all methods rely on pre-trained models and language-driven reasoning to bridge the gap between modalities.